A horizontal force of 99 N pushes a 11 kg block up a frictionless incline that makes an angle of 41° with the horizontal.

(a) What is the normal force that the incline exerts on the block?
                   =N

(b) What is the acceleration of the block? (Hint: assume up the block as the positive direction.)
                 =m/s2
Answer #1

Fsin41 41 deg Fcos41 masin41 mgmgco941 41 deg

Let's take the incline as the x-axis

Consider all forces in the y-direction

Fy=0

N - mgcos41 – Fsin41 = 0

N = mgcos41 + Fsin41

N = 11kg * 9.81m/s2 * cos41 +99 N * sin41

N = 11 * 9.81 * cos41 +99 * sin41

(a) ANSWER : V = 146.398

=============================

Consider the block

Net Force = Mass*acceleration

Fcos41 - mgsin 41 = ma

99 N * Cos41 - 11kg * 9.81m/s-sin41 = 11kg * a

99 * cos41 – 11 *9.81 sin41 = 11 *a

3.921 = 11*a

\frac{3.921}{11}=a

(b) ANSWER: a = 0.356m/s?

=============================
Answer #2

Draw the free body diagram of the block. Fcoso mgcos Fsin+mgcose/ 7.--- mg The normal reaction is expressed from the figure a

Substitute 9.8m/s2 for g, 11kg for m, 99 N for F, and 41° for e. N=(11kg) (9.8m/s) cos 41° +(99N)sin 41° = 146N (b) The net f